# Double Yangians and quantum vertex algebras, I

### v1

@@ 4. Vertex-operator Serre relation and iterate formula
Since
${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}=(z-w+\nu\hbar)a(z)b(w)=(z-w-\nu\hbar)b(w)a(z),$
we have ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}\in{\mathcal{E}}_{\hbar}^{(2)}(W)$, and
$a(z)b(w)=(z-w+\nu\hbar)^{-1}\ {\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}},\quad{\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}=-(w-z+\nu\hbar)b(w)a(z).$
Then
$z^{-1}\delta\!\left(\frac{w-\nu\hbar}{z}\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
$=\$$(z-w+\nu\hbar)^{-1}\ {\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}-(-w+z+\nu\hbar)^{-1}\ {\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
[…]
Applying $\operatorname{Res}_{z}$ and using Lemma 3.43, we get (4.25) as
${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z=w-\nu\hbar}=\operatorname{Res}_{z}Y_{{\mathcal{E}}}(a(w),z)b(w)=a(w)_{0}b(w).$
Next, we prove (4.26). Using definitions and delta-function substitution we have
$z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right)\!a(z_{1})\,{\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
[…]
and
$z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right)\left(\frac{z_{2}-z_{1}-2\nu\hbar}{z_{2}-z_{1}}\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}\;a(z_{1})$
[…]
$\times\left((z_{1}-w-\nu\hbar)(z_{2}-z_{1}-2\nu\hbar)(z_{2}-z_{1})^{-1}(z_{2}-w+\nu\hbar)a(z_{2})b(w)a(z_{1})\right)$
[…]
$\times\left((z_{1}-w+\nu\hbar)(z_{1}-z_{2}-2\nu\hbar)(z_{1}-z_{2})^{-1}(z_{2}-w+\nu\hbar)a(z_{1})a(z_{2})b(w)\right)$
[…]
Then we obtain
$z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right)\!a(z_{1}){\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}-z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right)\!\left(\!\frac{z_{2}-z_{1}-2\nu\hbar}{z_{2}-z_{1}}\!\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}a(z_{1})$
[…]
as desired. ∎
Now, we present the second characterization of the Serre relation.
@@
Let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module and let $a(z),b(z)\in{\mathcal{E}}_{\hbar}(W),\ \nu\in{\mathbb{C}}$ such that (4.21) and (4.22) hold. Then the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$ holds if and only if
(4.27) ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{1})a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w+\nu\hbar,z_{2}=w-\nu\hbar}=0.$
@@
By Proposition 4.5 (with $\hbar$ replaced by $\nu\hbar$), the very Serre relation is equivalent to
(4.28) $a(z_{1})\left(a(w)_{0}b(w)\right)=\left(a(w)_{0}b(w)\right)a(z_{1})\left(\frac{w-z_{1}-3\nu\hbar}{w-z_{1}-\nu\hbar}\right)\!,$
where (recall (4.25))
$a(w)_{0}b(w)=\operatorname{Res}_{x_{1}}\!\left(\!a(x_{1})b(w)-\left(\frac{w-x_{1}+\nu\hbar}{w-x_{1}-\nu\hbar}\right)b(w)a(x_{1})\!\right)\!={\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z=w-\nu\hbar}.$
Applying $\operatorname{Res}_{z_{2}}$ to (4.26) and using delta-function substitution, we get
(4.29) $a(z_{1})\left(a(w)_{0}b(w)\right)-\left(\frac{w-z_{1}-3\nu\hbar}{w-z_{1}-\nu\hbar}\right)\left(a(w)_{0}b(w)\right)a(z_{1})$
[…]
Then we see that (4.28) holds if and only if ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{1})a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w+\nu\hbar,z_{2}=w-\nu\hbar}=0$. ∎
Next, we consider the situation for $\hbar$-adic nonlocal vertex algebras and modules.
@@
Let $V$ be an $\hbar$-adic nonlocal vertex algebra with $a,b\in V$ such that
(4.30) $(z-w-2\hbar)Y(a,z)Y(a,w)=(z-w+2\hbar)Y(a,w)Y(a,z),$
(4.31) $(z-w+\hbar)Y(a,z)Y(b,w)=(z-w-\hbar)Y(b,w)Y(a,z).$
Then the order-$2$ Serre relation for $(Y(a,z_{1}),Y(a,z_{2}),Y(b,w))$ holds if and only if
(4.32) ${\rm Sing}_{z}Y(a,z)a_{0}b=0.$
@@
From Proposition 4.5 and Lemma 3.43, we see that the order-$2$ Serre relation for $(Y(a,z_{1}),Y(a,z_{2}),Y(b,w))$ holds if and only if
(4.33) $Y(a,z)Y(a_{0}b,w)-\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!Y(a_{0}b,w)Y(a,z)=0,$
which is equivalent to that
(4.34) $Y(a,z)Y(a_{0}b,w)\sim\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!Y(a_{0}b,w)Y(a,z)$
and ${\rm Sing}_{z}Y(a,z)a_{0}b=0$. With (4.31), by Lemma 4.3 we have
$x_{1}^{-1}\delta\left(\frac{w-\hbar}{x_{1}}\right)\!Y(a_{0}b,w)=Y(a,x_{1})Y(b,w)-\left(\frac{w-x_{1}+\hbar}{w-x_{1}-\hbar}\right)\!Y(b,w)Y(a,x_{1}).$
Then using commutation relations (4.30) and (4.31) we get
$(z-w+\hbar)(z-w-\hbar)Y(a,z)Y(a_{0}b,w)=(z-w+3\hbar)(z-w-\hbar)Y(a_{0}b,w)Y(a,z),$
which implies (4.34). Now, it follows. ∎
Proposition 4.10.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra and let $a,b\in V$. If
(4.35) $(z-w-2\hbar)Y(a,z)Y(a,w)=(z-w+2\hbar)Y(a,w)Y(a,z),$
(4.36) $(z-w+\hbar)Y(a,z)Y(b,w)=(z-w-\hbar)Y(b,w)Y(a,z),$
and if the order-$2$ Serre relation for $(Y(a,z_{1}),Y(a,z_{2}),Y(b,w))$ holds, then for any $V$-module $(W,Y_{W})$,
(4.37) $(z-w-2\hbar)Y_{W}(a,z)Y_{W}(a,w)=(z-w+2\hbar)Y_{W}(a,w)Y_{W}(a,z),$
(4.38) $(z-w+\hbar)Y_{W}(a,z)Y_{W}(b,w)=(z-w-\hbar)Y_{W}(b,w)Y_{W}(a,z),$
and the order-$2$ Serre relation for $(Y_{W}(a,z_{1}),Y_{W}(a,z_{2}),Y_{W}(b,w))$ holds. The converse is also true for any faithful $V$-module $(W,Y_{W})$.
Proof.
In view of Proposition 4.5, under the assumptions (4.35) and (4.36), the order-$2$ Serre relation for $(Y(a,z_{1}),Y(a,z_{2}),Y(b,w))$ holds if and only if
(4.39) $Y(a,z)Y(a_{0}b,w)=Y(a_{0}b,w)Y(a,z)\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!.$
For the same reason, under the assumption that (4.37) and (4.38) hold, the order-$2$ Serre relation for $(Y_{W}(a,z_{1}),Y_{W}(a,z_{2}),Y_{W}(b,w))$ holds if and only if
(4.40) $Y_{W}(a,z)Y_{W}(a_{0}b,w)=Y_{W}(a_{0}b,w)Y_{W}(a,z)\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!.$
Then it follows immediately from Proposition 3.22. ∎
@@
(4.41) $(x_{1}-x_{2}-2\hbar)a(x_{1})a(x_{2})=(x_{1}-x_{2}+2\hbar)a(x_{2})a(x_{1}),$
(4.42) $(x_{1}-x_{2}+\hbar)a(x_{1})b(x_{2})=(x_{1}-x_{2}-\hbar)b(x_{2})a(x_{1}).$
Assume that the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$ holds. Then
$(z_{1}-z_{2}-2\hbar)Y_{{\mathcal{E}}}(a(x),z_{1})Y_{{\mathcal{E}}}(a(x),z_{2})=(z_{1}-z_{2}+2\hbar)Y_{{\mathcal{E}}}(a(x),z_{2})Y_{{\mathcal{E}}}(a(x),z_{1}),$
$(z_{1}-z_{2}+\hbar)Y_{{\mathcal{E}}}(a(x),z_{1})Y_{{\mathcal{E}}}(b(x),z_{2})=(z_{1}-z_{2}-\hbar)Y_{{\mathcal{E}}}(b(x),z_{2})Y_{{\mathcal{E}}}(a(x),z_{1})$
on ${\langle}U{\rangle}$, and the order-$2$ Serre relation for $(Y_{{\mathcal{E}}}(a(x),z_{1}),Y_{{\mathcal{E}}}(a(x),z_{2}),Y_{{\mathcal{E}}}(b(x),w))$ holds.
@@ 5. Double Yangians and $\hbar$-adic quantum vertex algebras
In this section, for any simply-laced generalized Cartan matrix (GCM) $A$, we introduce an algebra $\widehat{\mathcal{DY}}(A)$, generalizing centrally extended double Yangians. As the main results, we give a new field presentation of $\widehat{\mathcal{DY}}(A)$ and a (tautological) construction of a universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ of any level $\ell\in{\mathbb{C}}$, and we prove that $\mathcal{V}_{A}(\ell)$ has a canonical $\hbar$-adic weak quantum vertex algebra structure with any (other) restricted $\widehat{\mathcal{DY}}(A)$-modules of level $\ell$ as modules.
@@ 5.1. Centrally extended double Yangian
Let $A=[a_{i,j}]_{i,j\in I}$ be a simply-laced GCM, which is fixed throughout this section. Note that $A$ is symmetric and
$a_{i,j}\in\{-1,0,2\}\quad\text{ for }i,j\in I.$
@@
$\textnormal{{(DY2)}}\quad H_{i}^{+}(z)H_{j}^{-}(w)=H_{j}^{-}(w)H_{i}^{+}(z)\frac{(z-w-a_{i,j}\hbar-\kappa\hbar)(z-w+a_{i,j}\hbar+\kappa\hbar)}{(z-w+a_{i,j}\hbar-\kappa\hbar)(z-w-a_{i,j}\hbar+\kappa\hbar)},$
$\textnormal{{(DY3)}}\quad H_{i}^{+}(z)X_{j}^{\pm}(w)=X_{j}^{\pm}(w)H_{i}^{+}(z)\left(\frac{z-w+a_{i,j}\hbar\pm{\frac{1}{2}}\kappa\hbar}{z-w-a_{i,j}\hbar\pm{\frac{1}{2}}\kappa\hbar}\right)^{\pm 1},$
$\textnormal{{(DY4)}}\quad H_{i}^{-}(z)X_{j}^{\pm}(w)=X_{j}^{\pm}(w)H_{i}^{-}(z)\left(\frac{w-z-a_{j,i}\hbar\pm{\frac{1}{2}}k\hbar}{w-z+a_{j,i}\hbar\pm{\frac{1}{2}}\kappa\hbar}\right)^{\pm 1},$
$\textnormal{{(DY5)}}\quad[X_{i}^{+}(z),X_{j}^{-}(w)]=\delta_{i,j}\frac{1}{2\hbar}\Bigg(\!H_{i}^{+}(w+{\frac{1}{2}}\kappa\hbar)z^{-1}\delta\!\left(\frac{w+\kappa\hbar}{z}\right)$
@@
$\textnormal{{(DY6)}}\quad(z-w\mp a_{i,j}\hbar)X_{i}^{\pm}(z)X_{j}^{\pm}(w)=(z-w\pm a_{i,j}\hbar)X_{j}^{\pm}(w)X_{i}^{\pm}(z),$
$\textnormal{{(DY7)}}\quad X_{i}^{\pm}(z)X_{j}^{\pm}(w)=X_{j}^{\pm}(w)X_{i}^{\pm}(z)\ \ \text{ if }a_{i,j}=0,$
$\textnormal{{(DY8)}}\quad\sum_{\sigma\in S_{2}}\Big(X_{i}^{\pm}(z_{\sigma(1)})X_{i}^{\pm}(z_{\sigma(2)})X_{j}^{\pm}(w)-2X_{i}^{\pm}(z_{\sigma(1)})X_{j}^{\pm}(w)X_{i}^{\pm}(z_{\sigma(2)})$
$\qquad\qquad+X_{j}^{\pm}(w)X_{i}^{\pm}(z_{\sigma(1)})X_{i}^{\pm}(z_{\sigma(2)})\Big)=0\quad\textnormal{{if }}a_{i,j}=-1.$
@@
Note that if $A$ is of finite type with ${\mathfrak{g}}={\mathfrak{g}}(A)$, this definition except (DY7) is the same as in [DK, page 17] (after a parameter-change $\hbar\rightarrow 2i\hbar$). On the other hand, if ${\mathfrak{g}}=\mathfrak{sl}_{n+1}$, $\widehat{\mathcal{DY}}(A)$ is isomorphic to the algebra $\mathcal{DY}_{\hbar}({\mathfrak{g}})$ defined in [I, Corollary 3.4].
@@
$[h_{i,\mathcal{Y}}(z),h_{j,\mathcal{Y}}(w)]=[a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}\left((z-w+\hbar\kappa)^{-2}-(w-z+\hbar\kappa)^{-2}\right)\!,$
$[h_{i,\mathcal{Y}}(z),x_{j,\mathcal{Y}}^{\pm}(w)]=\pm x_{j,\mathcal{Y}}^{\pm}(w)[a_{i,j}]_{q^{\partial_{w}}}\left((z-w+\hbar\kappa)^{-1}+(w-z+\hbar\kappa)^{-1}\right)\!,$
$(z-w-a_{i,j}\hbar)x_{i,\mathcal{Y}}^{\pm}(z)x_{j,\mathcal{Y}}^{\pm}(w)=(z-w+a_{i,j}\hbar)x_{j,\mathcal{Y}}^{\pm}(w)x_{i,\mathcal{Y}}^{\pm}(z),$
$x_{i,\mathcal{Y}}^{\pm}(z)x_{j,\mathcal{Y}}^{\pm}(w)=x_{j,\mathcal{Y}}^{\pm}(w)x_{i,\mathcal{Y}}^{\pm}(z)\ \ \text{ if }a_{i,j}=0,$
$x_{i,\mathcal{Y}}^{+}(z)x_{j,\mathcal{Y}}^{-}(w)-\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{+}(z)$
$\quad=\delta_{i,j}\frac{1}{2\hbar}\left(\!z^{-1}\delta\!\left(\frac{w}{z}\right)-C_{i}(w)z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\right)\!,$
and the order-$2$ Serre relations for $(x_{i,\mathcal{Y}}^{\pm}(z_{1}),x_{i,\mathcal{Y}}^{\pm}(z_{2}),x_{j,\mathcal{Y}}^{\pm}(w))$ hold if $a_{i,j}=-1$.
@@
$[h_{i,\mathcal{Y}}^{+}(z),h_{j,\mathcal{Y}}(w)]=[a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\hbar\kappa)^{-2},$
$[h_{i,\mathcal{Y}}^{-}(z),h_{j,\mathcal{Y}}(w)]=-[a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(w-z+\hbar\kappa)^{-2}.$
@@
(5.18) $[h_{i,\mathcal{Y}}^{+}(z),h_{j,\mathcal{Y}}^{-}(w)]=[a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\hbar\kappa)^{-2}$
for all $i,j\in I$ (recall that $a_{i,j}=a_{j,i}$).
@@
$\log\frac{(z-w-a_{i,j}\hbar-\kappa\hbar)(z-w+a_{i,j}\hbar+\kappa\hbar)}{(z-w+a_{i,j}\hbar-\kappa\hbar)(z-w-a_{i,j}\hbar+\kappa\hbar)}$
$=$ $\ -G(\partial_{z})G(\partial_{w})[a_{i,j}]_{q^{\partial_{z}}}[\kappa]_{q^{\partial_{z}}}(z-w)^{-2}.$
@@
$[\log H_{i}^{+}(z),\log H_{j}^{-}(w)]=-G(\partial_{z})G(\partial_{w})[a_{i,j}]_{q^{\partial_{z}}}[\kappa]_{q^{\partial_{z}}}(z-w)^{-2},$
@@
$=$ $\ [a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\kappa\hbar)^{-2}.$

### v2

@@ 4. Vertex-operator Serre relation and iterate formula
It is no harm to assume $\mu=1$. Applying Lemma 4.12 to Lemma 4.8, we get
$\sum_{\sigma\in S_{m}}\sum_{r=0}^{m}\binom{m}{r}(-1)^{r}a(z_{\sigma(1)})a(z_{\sigma(2)})\cdots a(z_{\sigma(r)})b(w)a(z_{\sigma(r+1)})\cdots a(z_{\sigma(m)})$
(4.30) $\quad=\sum_{\sigma\in S_{m}}a(w)_{0}^{m}b(w)\prod_{s=1}^{m}z_{\sigma(s)}^{-1}\delta\left(\frac{w-(\nu-2m+2s)\hbar}{z_{\sigma(s)}}\right).$
Then (4.28) implies the Serre-like relation for $(a(z),b(z),m)$. On the other hand, assume that the Serre-like relation holds for $(a(z),b(z),m)$. By multiplying $\prod_{s=1}^{m}\prod_{t\neq s}(z_{s}-w+(\nu-2m+2t)\hbar)$ and taking $\prod_{s=1}^{m}\operatorname{Res}_{z_{s}}$ on both hand sizes of (4.30), we get that
$\bigg(\prod_{s=1}^{m}\prod_{t\neq s}2(t-s)\hbar\bigg)a(w)_{0}^{m}b(w)=0.$
Therefore, $a(w)_{0}^{m}b(w)=0$. The equivalence between (4.28) and (4.29) follows immediate from (4.17). ∎
Next, we consider the situation with $\hbar$-adic nonlocal vertex algebras and modules.
@@
Let $V$ be an $\hbar$-adic nonlocal vertex algebra with $a,b\in V$ and $\mu\in{\mathbb{C}}^{\times}$, $\nu\in{\mathbb{N}}$ such that
(4.31) $(z-w-2\mu\hbar)Y(a,z)Y(a,w)=(z-w+2\mu\hbar)Y(a,w)Y(a,z),$
(4.32) $(z-w+\mu\nu\hbar)Y(a,z)Y(b,w)=(z-w-\mu\nu\hbar)Y(b,w)Y(a,z).$
Then the Serre-like relation holds for $(Y(a,z),Y(b,z),\nu+1)$ if and only if
(4.33) $a_{0}^{\nu+1}b=0.$
@@
From Theorem 4.13, we have that the Serre-like relation holds for
$(Y(a,z),Y(b,z),\nu+1)$
if and only if
$0=Y(a,z)_{0}^{\nu+1}Y(b,z)=Y(a_{0}^{\nu+1}b,z).$
Since $Y(\cdot,z)$ is injective, we complete the proof. ∎
@@
Let $V$ be an $\hbar$-adic nonlocal vertex algebra, let $a,b\in V$ and let $\mu\in{\mathbb{C}}^{\times}$, $\nu\in{\mathbb{N}}$. If
(4.34) $(z-w-2\mu\hbar)Y(a,z)Y(a,w)=(z-w+2\mu\hbar)Y(a,w)Y(a,z),$
(4.35) $(z-w+\mu\nu\hbar)Y(a,z)Y(b,w)=(z-w-\mu\nu\hbar)Y(b,w)Y(a,z),$
and if the Serre-like relation holds for $(Y(a,z),Y(b,z),\nu+1)$, then for any $V$-module $(W,Y_{W})$,
(4.36) $(z-w-2\mu\hbar)Y_{W}(a,z)Y_{W}(a,w)=(z-w+2\mu\hbar)Y_{W}(a,w)Y_{W}(a,z),$
(4.37) $(z-w+\mu\nu\hbar)Y_{W}(a,z)Y_{W}(b,w)=(z-w-\mu\nu\hbar)Y_{W}(b,w)Y_{W}(a,z),$
and the Serre-like relation holds for $(Y_{W}(a,z),Y_{W}(b,z),\nu+1)$. The converse is also true for any faithful $V$-module $(W,Y_{W})$.
@@
In view of Theorem 4.13, under the assumption (4.34) and (4.35), the Serre-like relation for $(Y(a,z),Y(b,z),\nu+1)$ holds if and only if
(4.38) $a_{0}^{\nu+1}b=0.$
For the same reason, under the assumption that (4.36) and (4.37) hold, the Serre-like relation for $(Y_{W}(a,z),Y_{W}(b,z),\nu+1)$ holds if and only if
(4.39) $0=Y_{W}(a,z)_{0}^{\nu+1}Y_{W}(b,z)=Y_{W}(a_{0}^{\nu+1}b,z),$
where the last equation follows from Lemma 3.38. Now it follows. ∎
@@
(4.40) $(x_{1}-x_{2}-2\mu\hbar)a(x_{1})a(x_{2})=(x_{1}-x_{2}+2\mu\hbar)a(x_{2})a(x_{1}),$
(4.41) $(x_{1}-x_{2}+\mu\nu\hbar)a(x_{1})b(x_{2})=(x_{1}-x_{2}-\mu\nu\hbar)b(x_{2})a(x_{1})$
for some $\mu\in{\mathbb{C}}^{\times}$, $\nu\in{\mathbb{N}}$. Assume that the Serre-like relation for $(a(z),b(z),\nu+1)$ holds. Then
$(z_{1}-z_{2}-2\mu\hbar)Y_{{\mathcal{E}}}(a(x),z_{1})Y_{{\mathcal{E}}}(a(x),z_{2})=(z_{1}-z_{2}+2\mu\hbar)Y_{{\mathcal{E}}}(a(x),z_{2})Y_{{\mathcal{E}}}(a(x),z_{1}),$
$(z_{1}-z_{2}+\mu\nu\hbar)Y_{{\mathcal{E}}}(a(x),z_{1})Y_{{\mathcal{E}}}(b(x),z_{2})=(z_{1}-z_{2}-\mu\nu\hbar)Y_{{\mathcal{E}}}(b(x),z_{2})Y_{{\mathcal{E}}}(a(x),z_{1})$
on ${\langle}U{\rangle}$, and the Serre-like relation holds for $(Y_{{\mathcal{E}}}(a(x),z),Y_{{\mathcal{E}}}(b(x),z),\nu+1)$.
@@ 5. Double Yangians and $\hbar$-adic quantum vertex algebras
In this section, for any symmetrizable generalized Cartan matrix (GCM) $A$, we introduce an algebra $\widehat{\mathcal{DY}}(A)$, generalizing centrally extended double Yangians. As the main results, we give a new field presentation of $\widehat{\mathcal{DY}}(A)$ and a (tautological) construction of a universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ of any level $\ell\in{\mathbb{C}}$, and we prove that $\mathcal{V}_{A}(\ell)$ has a canonical $\hbar$-adic weak quantum vertex algebra structure with any (other) restricted $\widehat{\mathcal{DY}}(A)$-modules of level $\ell$ as modules.
@@ 5.1. Centrally extended double Yangian
Let $A=[a_{i,j}]_{i,j\in I}$ be a symmetrizable GCM, which is fixed throughout this section. Then there are unique relatively prime positive integers $r_{i}$ ($i\in I$) such that $DA$ is symmetric with $D=\textnormal{{diag}}\{r_{i}\}_{i\in I}$.
@@
$\textnormal{{(DY2)}}\quad H_{i}^{+}(z)H_{j}^{-}(w)=H_{j}^{-}(w)H_{i}^{+}(z)\frac{(z-w-r_{i}a_{i,j}\hbar-\kappa\hbar)(z-w+r_{i}a_{i,j}\hbar+\kappa\hbar)}{(z-w+r_{i}a_{i,j}\hbar-\kappa\hbar)(z-w-r_{i}a_{i,j}\hbar+\kappa\hbar)},$
$\textnormal{{(DY3)}}\quad H_{i}^{+}(z)X_{j}^{\pm}(w)=X_{j}^{\pm}(w)H_{i}^{+}(z)\left(\frac{z-w+r_{i}a_{i,j}\hbar\pm{\frac{1}{2}}\kappa\hbar}{z-w-r_{i}a_{i,j}\hbar\pm{\frac{1}{2}}\kappa\hbar}\right)^{\pm 1},$
$\textnormal{{(DY4)}}\quad H_{i}^{-}(z)X_{j}^{\pm}(w)=X_{j}^{\pm}(w)H_{i}^{-}(z)\left(\frac{w-z-r_{i}a_{j,i}\hbar\pm{\frac{1}{2}}k\hbar}{w-z+r_{i}a_{j,i}\hbar\pm{\frac{1}{2}}\kappa\hbar}\right)^{\pm 1},$
$\textnormal{{(DY5)}}\quad[X_{i}^{+}(z),X_{j}^{-}(w)]=\delta_{i,j}\frac{1}{2r_{i}\hbar}\Bigg(\!H_{i}^{+}(w+{\frac{1}{2}}\kappa\hbar)z^{-1}\delta\!\left(\frac{w+\kappa\hbar}{z}\right)$
@@
$\textnormal{{(DY6)}}\quad(z-w\mp r_{i}a_{i,j}\hbar)X_{i}^{\pm}(z)X_{j}^{\pm}(w)=(z-w\pm r_{i}a_{i,j}\hbar)X_{j}^{\pm}(w)X_{i}^{\pm}(z),$
$\textnormal{{(DY7)}}\quad\sum_{\sigma\in S_{1-a_{i,j}}}\sum_{s=0}^{1-a_{i,j}}\binom{1-a_{i,j}}{s}(-1)^{s}X_{i}^{\pm}(z_{\sigma(1)})X_{i}^{\pm}(z_{\sigma(2)})\cdots X_{i}^{\pm}(z_{\sigma(s)})X_{j}^{\pm}(w)$
$\qquad\qquad\times X_{i}^{\pm}(z_{\sigma(s+1)})\cdots X_{i}^{\pm}(z_{\sigma(1-a_{i,j})})=0\quad\textnormal{{if }}a_{i,j}\leq 0.$
@@
Note that if $A$ is of simply-laced finite type with ${\mathfrak{g}}={\mathfrak{g}}(A)$, this definition except (DY7) for $a_{i,j}=0$ is the same as in [DK, page 37] (after a parameter-change $\hbar\rightarrow 2i\hbar$). On the other hand, if ${\mathfrak{g}}=\mathfrak{sl}_{n+1}$, $\widehat{\mathcal{DY}}(A)$ is isomorphic to the algebra $\mathcal{DY}_{\hbar}({\mathfrak{g}})$ defined in [I, Corollary 3.4].
@@
$[h_{i,\mathcal{Y}}(z),h_{j,\mathcal{Y}}(w)]=[r_{i}a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}\left((z-w+\hbar\kappa)^{-2}-(w-z+\hbar\kappa)^{-2}\right)\!,$
$[h_{i,\mathcal{Y}}(z),x_{j,\mathcal{Y}}^{\pm}(w)]=\pm x_{j,\mathcal{Y}}^{\pm}(w)[r_{i}a_{i,j}]_{q^{\partial_{w}}}\left((z-w+\hbar\kappa)^{-1}+(w-z+\hbar\kappa)^{-1}\right)\!,$
$(z-w-r_{i}a_{i,j}\hbar)x_{i,\mathcal{Y}}^{\pm}(z)x_{j,\mathcal{Y}}^{\pm}(w)=(z-w+r_{i}a_{i,j}\hbar)x_{j,\mathcal{Y}}^{\pm}(w)x_{i,\mathcal{Y}}^{\pm}(z),$
$x_{i,\mathcal{Y}}^{+}(z)x_{j,\mathcal{Y}}^{-}(w)-\left(\frac{w-z+r_{i}a_{i,j}\hbar}{w-z-r_{i}a_{i,j}\hbar}\right)x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{+}(z)$
$\quad=\delta_{i,j}\frac{1}{2r_{i}\hbar}\left(\!z^{-1}\delta\!\left(\frac{w}{z}\right)-C_{i}(w)z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\right)\!,$
and the Serre-like relations for $(x_{i,\mathcal{Y}}^{\pm}(z),x_{j,\mathcal{Y}}^{\pm}(z),1-a_{i,j})$ hold if $a_{i,j}\leq 0$.
@@
$[h_{i,\mathcal{Y}}^{+}(z),h_{j,\mathcal{Y}}(w)]=[r_{i}a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\hbar\kappa)^{-2},$
$[h_{i,\mathcal{Y}}^{-}(z),h_{j,\mathcal{Y}}(w)]=-[r_{i}a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(w-z+\hbar\kappa)^{-2}.$
@@
(5.18) $[h_{i,\mathcal{Y}}^{+}(z),h_{j,\mathcal{Y}}^{-}(w)]=[r_{i}a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\hbar\kappa)^{-2}$
for all $i,j\in I$ (recall that $r_{i}a_{i,j}=r_{j}a_{j,i}$).
@@
$\log\frac{(z-w-r_{i}a_{i,j}\hbar-\kappa\hbar)(z-w+r_{i}a_{i,j}\hbar+\kappa\hbar)}{(z-w+r_{i}a_{i,j}\hbar-\kappa\hbar)(z-w-r_{i}a_{i,j}\hbar+\kappa\hbar)}$
$=$ $\ -G(\partial_{z})G(\partial_{w})[r_{i}a_{i,j}]_{q^{\partial_{z}}}[\kappa]_{q^{\partial_{z}}}(z-w)^{-2}.$
@@
$[\log H_{i}^{+}(z),\log H_{j}^{-}(w)]=-G(\partial_{z})G(\partial_{w})[r_{i}a_{i,j}]_{q^{\partial_{z}}}[\kappa]_{q^{\partial_{z}}}(z-w)^{-2},$
@@
$=$ $\ [r_{i}a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\kappa\hbar)^{-2}.$